OpenStudy (anonymous):

Find Devivaties with respect to x: 1). y=sin(cos(x+3)); 2). tan(x+y)=ln x+ 5y.

OpenStudy (cwrw238):

1. - use the chain rule dy/dx = cos (cos(x+3) * sin (x + 3)

OpenStudy (anonymous):

i thought so but then i was thinking product rule...this is a review question for my calc 2 class

OpenStudy (cwrw238):

chain rule: - look for a 'function within a function'

OpenStudy (anonymous):

1). y=sin(cos(x+3)); 2). tan(x+y)=ln x+ 5y. 1:y = sin(cos(x+3)) y' = cos(cos(x+3))(-sin(x+3))(1) y' = {-sin(x+3)}{cos(cos(x+3))} 2. Tan(x+y) = lnx+5y Sec^2(x+y)(1+y') = 1/x + 5yy' y' = {((1/x))-(Sec^2(x+y))}/(Sec^2(x+y) - 5y) My calculus is rough right now. 2 im not sure but 1 im sure

OpenStudy (cwrw238):

product rule is used for product of 2 or more functions

OpenStudy (cwrw238):

oh yes - i made a mistake on no. 1:- the derivative of cos is - sin franciscanmpnk is correct

OpenStudy (anonymous):

yeaa i caught that ...his number 2 seems correct as well ...Thanks to the both of you!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

for no 1: {-sin(x+3)}*{cos(cos(x+3)}

OpenStudy (anonymous):

yes your correct...how about number 2...need a second opinion

OpenStudy (cwrw238):

no 2 is implicit differentiation where y is taken to be a function of x tan(x+y)=ln x+ 5y. use chain rule for tan(x + y) = sec^2(x + y) * (1 + y') ln x + 5y = 1/x + 5yy' sec^2(x + y) * (1 + y') = 1/x + 5yy' now we need to make y' the subject

OpenStudy (anonymous):

make sense

OpenStudy (anonymous):

makes sense i mean

OpenStudy (anonymous):

yes cw, that is that is just what i would do.

OpenStudy (cwrw238):

sec^2(x + y) + y' sec^2(x + y) = 1/x + 5yy' y' sec^2(x + y) - 5yy' = 1/x - sec^2(x + y) y' = [ 1/x - sec^2(x + y) ] / [sec^2(x + y) - 5y]

OpenStudy (cwrw238):

- same as franciscanmonk got

OpenStudy (anonymous):

thaks cw...you explain things very clear

OpenStudy (cwrw238):

ty tvvv1012 but there is a mistake in the above the derivative of 5y is NOT 5yy' but 5y'

OpenStudy (cwrw238):

so both francis and i are wrong

OpenStudy (cwrw238):

y' sec^2(x + y) - 5y' = 1/x - sec^2(x + y) y' = [ 1/x - sec^2(x + y) ] / [sec^2(x + y) - 5]

OpenStudy (anonymous):

haha sorry about no.2

OpenStudy (cwrw238):

i made the same mistake at first - hey - shows we're human!!!